OpenStudy (anonymous):

@DemolisionWolf Sorry to bother you again, do you think you can help me with one more though?

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

Part A i suppose is where you need some direction?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so to factor -16t^2 - 64t + 512 we need to find the factors of each coefficients, which are 16, 64, and 512 looking at it, i can see that 8 * 2 = 16 8 * 8 = 64 8 * 8 * 8 = 512 so, lets factor out a 8 out of all of it 8 ( -8t^2 - 8t + 64) make sense so far?

OpenStudy (anonymous):

8 ( -2t^2 - 8t + 64) sorry

OpenStudy (anonymous):

now, that it's simplier, I can see that each coefficeint has a factor of 2, so i'll pull a 2 out of each term (8)(2)(-t^2 - 4t + 32)

OpenStudy (anonymous):

(8)(2)(-t^2 - 4t + 32) becomes (16)(-t^2 - 4t + 32) I can't have a -t^2, so i'll take the negative out and put it with the 16 (-16)(t^2 + 4t - 32) now it's time to Factor, or FOIL in reverse, i'm going to ignore the (16) in front for a bit (t^2 + 4t - 32) i know this much (t )(t ) now I need to figure out factors of 32 32 = 2*16 32 = 4*8 since the 4t term has a 4, and the 4*8=32 the 4 and 8 are different by 4, i'm going to pick the factors of 4*8 (t 4)(t 8) now i need to put in the correct signs i put the + with the 8 because my 4t term is positive (t - 4)(t + 8) so, those are my factors for t!

OpenStudy (anonymous):

and u logged off.... >.<

OpenStudy (anonymous):

Sorry, I had to take a break from all the math. Thank you so much for your help.

OpenStudy (anonymous):

^_^ your questions are long and quite dificult..

OpenStudy (anonymous):

Yeah I know, there was 5 of them. That's why I needed to take a break.

OpenStudy (anonymous):

haha your questions require ton's of algebra steps... like methods from like 5 chapters, haha

OpenStudy (anonymous):

Haha, yeah, tell me about it. It gets very frustrating most of the time.

OpenStudy (anonymous):

what college class is it, physics?

OpenStudy (anonymous):

Actually, it's high school Algebra. Lmao

OpenStudy (anonymous):

oh rad... haha